You have a population with three alleles: G_1, has
Answer #1

a).

Given that the frequency of allele, G1 = 0.1, G2 = 0.7, G3 = 0.2.

The Hardy-Weinberg equation for three alleles p, q and r is, p2 (G1G1)+ 2pq (G1G2) +q2 (G2G2) + 2pr (G1G3)+ 2qr (G2G3) + r2 (G3G3) = 1

Now, the frequency of each genotype, G1G1, p2 = 0.1*0.1 = 0.01

The frequency of G1G2, or 2pq = 2*0.1* 0.7 = 0.14

The frequency of G2G2 or q2 = 0.7*0.7 = 0.49

The frequency of G1G3 or 2pr = 2*0.1*0.2 = 0.04

The frequency of G2G3 or 2qr = 2* 0.7* 0.2 = 0.28

The frequency of G3G3 or q2 = 0.2*0.2 = 0.04

The frequencies of all genotypes = 0.01+0.14+0.49+0.04+0.28+0.04 = 1


b). If the population is in Hardy-Weinberg equilibrium, all the genotypic frequencies must be equal to 1. The given population is in Hardy-Weinberg equilibrium, as it is equal to 1.